OpenStudy (scoobydo):

Which of the following sets shows all the numbers from the set {2, 3, 4, 5} that are part of the solution to the inequality 4n + 6 > 22

OpenStudy (scoobydo):

metal and fan

OpenStudy (anonymous):

Okay, so we can begin by going down the line and just seeing what does and does not work. 2 x 4 = 8. 8 + 6 = 14. 4 x 3 = 12. 12 + 6 = 18. 4 x 4 = 16. 16 + 6 = 22. All we have left is 5, and it just so happens 5 is the answer! If I helped, a medal and a fan would be appreciated :D